ICM 1998, Berlin, Aug. 18–27

Abstracts of Plenary and Invited Lectures

Section:

2. Algebra

1991 MS Classification: William M. Kantor, Department of Mathematics University of Oregon Eugene, OR 97403-1222 USA Simple Groups in Computational Group Theory

Given an arbitrary subset S of the symmetric group S_{n}, can one efficiently find properties of G = ⟨S⟩, such as |G|, the derived series, a composition series, Sylow subgroups … ? I will briefly survey recent approaches to such questions, emphasizing the use of simple groups and of probabilistic methods for devising algorithms.

ICM 1998 Berlin · Plenary and Invited Lectures